# ALMA-IMF XV: NH kinematic analysis on the intermediate protocluster G353.41 Key Words.: stars: formation – ISM: clouds – ISM: kinematics and dynamics – ISM: molecules1 Introduction

### v1

@@ Acknowledgements.
This paper makes use of the following ALMA data: ADS/JAO.ALMA#2017.1.01355.L. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. We thank Elena Redaelli and Diego R. Matus Carrillo for very helpful discussions. R.A. gratefully acknowledges support from ANID Beca Doctorado Nacional 21200897. A.S. gratefully acknowledges support by the Fondecyt Regular (project code 1220610) and ANID BASAL project FB210003. F.L. acknowledges the support of the Marie Curie Action of the European Union (project MagiKStar, Grant agreement number 841276) F.M. acknowledges support from the French Agence Nationale de la Recherche (ANR) under reference ANR-20-CE31-009, of the Programme National de Physique Stellaire and Physique et Chimie du Milieu Interstellaire (PNPS and PCMI) of CNRS/INSU (with INC/INP/IN2P3). R.G.M. and D.D.G. acknowledge support from UNAM-PAPIIT project IN108822 and from CONACyT Ciencia de Frontera project ID 86372. F.M., F.L., and N.C. acknowledge support from the European Research Council (ERC) via the ERC Synergy Grant ECOGAL (grant 855130). N.C. acknowledges funding from the ERC under the European Union’s Horizon 2020 research. P.S. was partially supported by a Grant-in-Aid for Scientific Research (KAKENHI Number JP22H01271 and JP23H01221) of JSPS. M.B. is a postdoctoral fellow in the University of Virginia’s VICO collaboration and is funded by grants from the NASA Astrophysics Theory Program (grant number 80NSSC18K0558) and the NSF Astronomy & Astrophysics program (grant number 2206516). T.Cs. has received financial support from the French State in the framework of the IdEx Université de Bordeaux Investments for the future Program. S.D.R. acknowledges the funding and support of ANID-Subdirección de Capital Humano Magíster/Nacional/2021-22211000. T.B. acknowledges the support from S. N. Bose National Centre for Basic Sciences under the Department of Science and Technology, Govt. of India. G.B. acknowledges financial support from the grants PID2020-117710GB-I00 and CEX2019-000918 funded by MCIN/AEI/10.13039/501100011033. A.K. and L.B. gratefully acknowledge support from ANID BASAL project FB210003. F.O. acknowledge the support of the Ministry of Science and Technology of Taiwan, projects No. 109-2112-M-007-008-, 110-2112-M-007-023-, and 110-2112-M-007-034-.References
@@ Appendix A Filamentary identification with FilFinder
For this approach we use the moment 0 map of the extracted NH isolated components that present a SNR 5. To estimate the moment 0 map we used the moment task from the SpectralCube Python package, within the velocity range of km s to 0 km s. As part of the pre-processing of the moment 0 map before the filamentary detection, we decrease the contrast in the image by using the preprocess_image task and its argument flatten_percent set to 90. Now, in order to indicate to FilFinder the area we to identify filaments use the subtask create_mask with the following parameters: glob_thresh: 4.5 K km s, size_thresh: 0.25 pc, smooth_size: 0.12 pc, border_masking: False, fill_hole_size: 0.013 pc. The resulting mask is presented in Fig. A.1 with a white contour.
Then, we obtain the skeletons of the mask by using medskel. The derived structures are presented with red and black lines in Fig. A.1. Given we are only interested in the large scale filaments, we use analyze_skeletons in order to “prune” the small scale structures. For this pruning we use branch_thresh: 0.3 pc, prune_criteria: ’length’, max_prune_iter: 0. This approach results in removing the small filaments (red lines in Fig. A.1) from the original skeleton and to obtain the main filamentary structure in G353 (black lines in Fig. A.1).
@@ Appendix B Examples of the isolated components fitting
Appendix C DCN & NH derived core velocity
@@ Appendix D V-shaped structures
In § 5.3 we characterized the most prominent V-shaped structure we detect in Fig. 9. We repeat this process for other eight different V-shaped structures, including the linear fits to the velocity gradients. In Fig. D.1 we indicate their location in PV space with arrows and present individual close-up for each V-shape in Fig. D.2. In Table D.1 we list their VGs, timescales, H masses, and mass accretion rates.
@@
In V-shapes “G” and “H” we see the same 1.3 mm cores with NH velocities. These V-shapes are not the same distribution. They are overlapped in PV space and spatially separated by .
We improve the clarity of V-shape “B” by rotating the data in PP space by 80 counter-clockwise. We apply this process for V-shapes “E”, “G”, and “H” with an angle of 33 clockwise.
@@ Appendix E SiO Intensity-weighted position-velocity diagram
To create the SiO intensity-weighted PV diagram, first, we remove most of the noisy spectra by considering data with . Then, we estimate the integrated intensity and velocity centroid at each pixel. We find improvements in our cleaning by using only spectra with integrated intensity K km s. Using the coordinate, integrated-intensity, and velocity centroid of each spectrum, we create the SiO intensity-weighted PV diagrams we show in E.1.
Core | RA | DEC | F | F | PA | Mass | V | Type |
Number | [ ] | [ ] | [″] | [″] | [ ] | [M ] | [km s] | |
2 | 262.6165032 | -34.6955865 | 0.065 | DCN, Single | ||||
3 | 262.6184156 | -34.6965240 | 0.040 | DCN, Single | ||||
4 | 262.6103159 | -34.6932659 | 0.155 | DCN, Single | ||||
5 | 262.6101515 | -34.6960014 | 0.087 | DCN, Single | ||||
6 | 262.6049155 | -34.6934384 | 0.126 | DCN, Single | ||||
7 | 262.6137738 | -34.6947298 | DCN & NH | |||||
8 | 262.6039531 | -34.6936374 | ||||||
9 | 262.6192359 | -34.6903650 | DCN & NH | |||||
11 | 262.6243189 | -34.6880780 | DCN & NH | |||||
12 | 262.6072148 | -34.6969795 | 0.019 | DCN, Single | ||||
13 | 262.6078228 | -34.6996836 | ||||||
14 | 262.6147937 | -34.6946762 | 0.090 | DCN, Single | ||||
15 | 262.6107433 | -34.6964412 | 0.087 | DCN, Complex | ||||
16 | 262.6215941 | -34.6989408 | ||||||
17 | 262.5954514 | -34.6916168 | ||||||
18 | 262.5927434 | -34.7052494 | ||||||
19 | 262.6064012 | -34.7019756 | ||||||
20 | 262.6111096 | -34.6932787 | ||||||
21 | 262.6131910 | -34.6939495 | 0.063 | DCN, Single | ||||
22 | 262.6118441 | -34.6946150 | DCN & NH | |||||
23 | 262.6028175 | -34.6925438 | ||||||
24 | 262.6198349 | -34.6960383 | 0.089 | DCN, Single | ||||
25 | 262.6155222 | -34.6952591 | DCN & NH | |||||
26 | 262.6143434 | -34.6917027 | ||||||
27 | 262.6000802 | -34.6910324 | ||||||
28 | 262.6253977 | -34.6999713 | ||||||
29 | 262.6133074 | -34.6919187 | ||||||
30 | 262.6114686 | -34.6962602 | 0.074 | DCN, Single | ||||
31 | 262.6096651 | -34.6925680 | ||||||
32 | 262.6094126 | -34.6910985 | ||||||
33 | 262.6287106 | -34.6862068 | ||||||
34 | 262.5982914 | -34.6919006 | 0.080 | DCN, Single | ||||
35 | 262.6142011 | -34.6940134 | DCN & NH | |||||
36 | 262.6202758 | -34.7001995 | ||||||
37 | 262.6010398 | -34.6950114 | 0.053 | DCN, Single | ||||
38 | 262.5971034 | -34.6920396 | ||||||
39 | 262.6054437 | -34.6963773 | DCN & NH | |||||
40 | 262.5917454 | -34.6897316 | ||||||
41 | 262.6095777 | -34.6983259 | ||||||
42 | 262.5975992 | -34.6876666 | ||||||
43 | 262.6035166 | -34.6966807 | 0.062 | DCN, Single | ||||
44 | 262.6030115 | -34.6956424 | 0.091 | DCN, Single | ||||
45 | 262.6143008 | -34.6909376 | ||||||
46 | 262.6187648 | -34.6912377 | DCN & NH | |||||
47 | 262.6178453 | -34.6919943 |
@@
V-shape ID | (H) | Upper / lower VG | Mean VG | Upper / lower | (H) | |||
[ ] | [ ] | [M] | [km s pc] | [km s pc] | [kyr] | [kyr] | [10 M yr] | |
@@ Appendix F G353 power law density profile
To determine the value of , we integrate this expression in a sphere (Eq. F.2), with pc. Based of different testings we set the total mass of the sphere to 150 M. We define which corresponds to the pixel size of the NH data at a distance of 2 kpc.
where ) = 150 M, pc, and = 5.65, from Eq. F.4, we obtain:

### v2

@@ Acknowledgements.
The authors thank the referee for helpful comments that improved the text. This paper makes use of the following ALMA data: ADS/JAO.ALMA#2017.1.01355.L. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. We thank Elena Redaelli, Diego R. Matus Carrillo, and Vineet Rawat for very helpful discussions. R.A. gratefully acknowledges support from ANID Beca Doctorado Nacional 21200897. A.S. gratefully acknowledges support by the Fondecyt Regular (project code 1220610) and ANID BASAL project FB210003. F.L. acknowledges the support of the Marie Curie Action of the European Union (project MagiKStar, Grant agreement number 841276) F.M. acknowledges support from the French Agence Nationale de la Recherche (ANR) under reference ANR-20-CE31-009, of the Programme National de Physique Stellaire and Physique et Chimie du Milieu Interstellaire (PNPS and PCMI) of CNRS/INSU (with INC/INP/IN2P3). R.G.M. and D.D.G. acknowledge support from UNAM-PAPIIT project IN108822 and from CONACyT Ciencia de Frontera project ID 86372. F.M., F.L., and N.C. acknowledge support from the European Research Council (ERC) via the ERC Synergy Grant ECOGAL (grant 855130). N.C. acknowledges funding from the ERC under the European Union’s Horizon 2020 research. P.S. was partially supported by a Grant-in-Aid for Scientific Research (KAKENHI Number JP22H01271 and JP23H01221) of JSPS. M.B. is a postdoctoral fellow in the University of Virginia’s VICO collaboration and is funded by grants from the NASA Astrophysics Theory Program (grant number 80NSSC18K0558) and the NSF Astronomy & Astrophysics program (grant number 2206516). A.G. acknowledges support from the NSF under grants AAG 2008101 and CAREER 2142300. T.Cs. has received financial support from the French State in the framework of the IdEx Université de Bordeaux Investments for the future Program. S.D.R. acknowledges the funding and support of ANID-Subdirección de Capital Humano Magíster/Nacional/2021-22211000. T.B. acknowledges the support from S. N. Bose National Centre for Basic Sciences under the Department of Science and Technology, Govt. of India. G.B. acknowledges financial support from the grants PID2020-117710GB-I00 and CEX2019-000918 funded by MCIN/AEI/10.13039/501100011033. A.K. and L.B. gratefully acknowledge support from ANID BASAL project FB210003. F.O. acknowledge the support of the Ministry of Science and Technology of Taiwan, projects No. 109-2112-M-007-008-, 110-2112-M-007-023-, and 110-2112-M-007-034-.References
@@
- Galván-Madrid et al. (2013) Galván-Madrid, R., Liu, H. B., Zhang, Z. Y., et al. 2013, ApJ, 779, 121
- Galván-Madrid et al. (2010) Galván-Madrid, R., Zhang, Q., Keto, E., et al. 2010, ApJ, 725, 17
@@
- Henshaw et al. (2020) Henshaw, J. D., Kruijssen, J. M. D., Longmore, S. N., et al. 2020, Nature Astronomy, 4, 1064
@@ Appendix A Filamentary identification with FilFinder
For this approach we use the moment 0 map of the extracted N2H+ isolated components that present a SNR 5. To estimate the moment 0 map we used the moment task from the SpectralCube Python package, within the velocity range of km s-1 to 0 km s-1. As part of the pre-processing of the moment 0 map before the filamentary detection, we decrease the contrast in the image by using the preprocess_image task and its argument flatten_percent set to 90. Now, in order to indicate to FilFinder the area we to identify filaments use the subtask create_mask with the following parameters: glob_thresh: 4.5 K km s-1, size_thresh: 0.25 pc2, smooth_size: 0.12 pc, border_masking: False, fill_hole_size: 0.013 pc2. The resulting mask is presented in Fig. A.1 with a white contour.
Then, we obtain the skeletons of the mask by using medskel. The derived structures are presented with red and green lines in Fig. A.1. Given we are only interested in the large scale filaments, we use analyze_skeletons in order to “prune” the small scale structures. For this pruning we use branch_thresh: 0.3 pc, prune_criteria: ’length’, max_prune_iter: 0. This approach results in removing the small filaments (red lines in Fig. A.1) from the original skeleton and to obtain the main filamentary structure in G353 (green lines in Fig. A.1).
@@ Appendix B Examples of the isolated components fitting
Appendix C DCN & N2H+ derived core velocity
@@ Appendix D V-shaped structures
In § 5.3 we characterized the most prominent V-shaped structure we detect in Fig. 9. We repeat this process for other eight different V-shaped structures, including the linear fits to the velocity gradients. In Fig. D.1 we indicate the V-shapes location in PV space with dark points, arrows, and their IDs. In Fig. D.2 we present individual close-ups for each V-shape. In Table D.1 we list their VGs, timescales, H2 masses, and mass accretion rates.
@@
In V-shapes “G” and “H” we see the same 1.3 mm cores with N2H+ velocities. These V-shapes are not the same distribution. They are overlapped in PV space and spatially separated by .
We improve the clarity of V-shape “B” by rotating the data in PP space by 80∘ counter-clockwise. We apply this process for V-shapes “E”, “G”, and “H” with an angle of 33∘ clockwise.
@@ Appendix E SiO Intensity-weighted position-velocity diagram
To create the SiO intensity-weighted PV diagram, first, we remove most of the noisy spectra by considering data with . Then, we estimate the integrated intensity and velocity centroid at each pixel. We find improvements in our cleaning by using only spectra with integrated intensity K km s-1. Using the coordinate, integrated-intensity, and velocity centroid of each spectrum, we create the SiO intensity-weighted PV diagrams we show in E.1.
Core | RA | DEC | FA | FB | PA | Mass | VLSR | Type | Number of | ||
Number | [∘ ] | [∘ ] | [″] | [″] | [∘ ] | [M⊙ ] | [km s-1] | N2H+ components | |||
2 | 262.6165032 | -34.6955865 | 0.065 | DCN, Single | |||||||
3 | 262.6184156 | -34.6965240 | 0.040 | DCN, Single | |||||||
4 | 262.6103159 | -34.6932659 | 0.155 | DCN, Single | |||||||
5 | 262.6101515 | -34.6960014 | 0.087 | DCN, Single | |||||||
6 | 262.6049155 | -34.6934384 | 0.126 | DCN, Single | |||||||
7 | 262.6137738 | -34.6947298 | DCN & N2H+ | ||||||||
8 | 262.6039531 | -34.6936374 | — | — | — | ||||||
9 | 262.6192359 | -34.6903650 | DCN & N2H+ | ||||||||
11 | 262.6243189 | -34.6880780 | DCN & N2H+ | ||||||||
12 | 262.6072148 | -34.6969795 | 0.019 | DCN, Single | |||||||
13 | 262.6078228 | -34.6996836 | — | — | — | ||||||
14 | 262.6147937 | -34.6946762 | 0.090 | DCN, Single | |||||||
15 | 262.6107433 | -34.6964412 | 0.087 | DCN, Complex | — | ||||||
16 | 262.6215941 | -34.6989408 | — | — | — | ||||||
17 | 262.5954514 | -34.6916168 | — | — | — | ||||||
18 | 262.5927434 | -34.7052494 | — | — | — | ||||||
19 | 262.6064012 | -34.7019756 | — | — | — | ||||||
20 | 262.6111096 | -34.6932787 | — | — | — | ||||||
21 | 262.6131910 | -34.6939495 | 0.063 | DCN, Single | |||||||
22 | 262.6118441 | -34.6946150 | DCN & N2H+ | ||||||||
23 | 262.6028175 | -34.6925438 | — | — | — | ||||||
24 | 262.6198349 | -34.6960383 | 0.089 | DCN, Single | |||||||
25 | 262.6155222 | -34.6952591 | DCN & N2H+ | ||||||||
26 | 262.6143434 | -34.6917027 | — | — | — | ||||||
27 | 262.6000802 | -34.6910324 | — | — | — | ||||||
28 | 262.6253977 | -34.6999713 | — | — | — | ||||||
29 | 262.6133074 | -34.6919187 | — | — | — | ||||||
30 | 262.6114686 | -34.6962602 | 0.074 | DCN, Single | |||||||
31 | 262.6096651 | -34.6925680 | — | — | — | ||||||
32 | 262.6094126 | -34.6910985 | — | — | — | ||||||
33 | 262.6287106 | -34.6862068 | — | — | — | ||||||
34 | 262.5982914 | -34.6919006 | 0.080 | DCN, Single | |||||||
35 | 262.6142011 | -34.6940134 | DCN & N2H+ | ||||||||
36 | 262.6202758 | -34.7001995 | — | — | — | ||||||
37 | 262.6010398 | -34.6950114 | 0.053 | DCN, Single | |||||||
38 | 262.5971034 | -34.6920396 | — | — | — | ||||||
39 | 262.6054437 | -34.6963773 | DCN & N2H+ | ||||||||
40 | 262.5917454 | -34.6897316 | — | — | — | ||||||
41 | 262.6095777 | -34.6983259 | — | — | — | ||||||
42 | 262.5975992 | -34.6876666 | — | — | — | ||||||
43 | 262.6035166 | -34.6966807 | 0.062 | DCN, Single | |||||||
44 | 262.6030115 | -34.6956424 | 0.091 | DCN, Single | |||||||
45 | 262.6143008 | -34.6909376 | — | — | — | ||||||
46 | 262.6187648 | -34.6912377 | DCN & N2H+ | ||||||||
47 | 262.6178453 | -34.6919943 | — | — | — |
@@
V-shape ID | (H2) | Upper / lower VG | Mean VG | Upper / lower | (H2) | |||
[∘ ] | [∘ ] | [M⊙] | [km s-1 pc-1] | [km s-1 pc-1] | [kyr] | [kyr] | [10-4 M⊙ yr-1] | |
@@ Appendix F G353 power law density profile
To determine the value of , we integrate this expression in a sphere (Eq. F.2), with pc. Based on different tests, probing total masses from M⊙ and values from , we set the total mass of the sphere to 150 M⊙. We define which corresponds to the pixel size of the N2H+ data at a distance of 2 kpc.
where ) = 150 M⊙, pc, and = 5.65, from Eq. F.4, we obtain: